calculate the energy expended per second by a person at 3m/sec (9-min.mile) with a muscle efficiency of 20% .assume that the mass m=10kg and the step length is 1m
1
Expert's answer
2019-01-04T10:20:09-0500

A human organism produces energy from food and breathing. This energy wastes for heat, mechanical motion (20% of all energy according to the condition) and other important inner processes like metabolism, thinking and solving problems.

These 20% of energy are used to support motion of 10 kg-body with the speed 3 m/s - to give you the kinetic energy:

"0.2E=mv^2\/2,""E=5mv^2\/2=5*10*3^2\/2=225"

J/s, or 53.8 calories per second, or 193.68 kcal per hour.

This energy is produced with one step. In 3 meters there are 3 steps. The more steps per second you do - the more energy you need to support this motion. Multiply the result by 3:

"E=3*225=675"

J/s or 161.4 cal/s or 581 kcal/h